MarkSucks:

Help pls :0 A jet made a trip to Jakarta and back. The trip there took 11 hours and the trip back took nine hours. It averaged 56 mph faster on the return trip than on the outbound trip. Find the jet's average speed on the outbound trip

Vocaloid:

in general, distance = speed * time we use this to write two equations: one for the outbound trip, one for the return trip we want to know the speed of the outbound trip, so let's let x = speed of outbound trip outbound trip: distance = (x)(11) return trip: the return speed is 56 hours than the outbound trip, so the speed is x + 56 and the return equation is: distance = (x+56)(9) now, since this is a trip back-and-forth from the same place, both distances are equal. thus: (x)(11) = (x+56)(9) and solve for x